OpenStudy (anonymous):

A cone has a volume of 6 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? A. 12 in3 B. 18 in3 C. 24 in3 D. 36 in3

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (mertsj):

Now let's think for a moment. What is the formula for the volume of a cone?

OpenStudy (anonymous):

1/3 pi(r)^2

OpenStudy (mertsj):

Almost.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[Vcylinder =πr^2h\]

OpenStudy (anonymous):

oh "h"!!!!

OpenStudy (mertsj):

\[V _{_{c}}=\frac{1}{3}\pi r^2h\]

OpenStudy (mertsj):

What is the formula for the volume of a cylinder?

OpenStudy (anonymous):

pi(r)^2(h)

OpenStudy (mertsj):

Yes. How do the volumes compare?

OpenStudy (anonymous):

The cone has 1/3

OpenStudy (mertsj):

Or another way to say that is "the volume of the cylinder is 3 times the volume of the cone."

OpenStudy (anonymous):

oh ok

OpenStudy (mertsj):

Do you understand? Of course the radii and heights must be the same.

OpenStudy (anonymous):

yes, but I am still confused how to figure out the volume of the cylinder so the cone can fit in it.

OpenStudy (anonymous):

The question only gives me the volume of the cone, do I just multiply that number by 3? So 6 * 3 = 18?

OpenStudy (anonymous):

Are you there? @Mertsj

OpenStudy (mertsj):

yes and yes

OpenStudy (anonymous):

ok so 18 is the answer?

OpenStudy (anonymous):

Can I ask you another question?

OpenStudy (mertsj):

yes. The idea is that if you have a cylinder and cone of equal radii and height, then 3 of the cones will fit in the cylinder.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hold on one sec, I have one more question

OpenStudy (mertsj):

Are you going to post it or am I supposed to guess what it is?

OpenStudy (anonymous):

Betty closes the nozzle of the funnel and fills it completely with a liquid. She then opens the nozzle. If the liquid drips at the rate of 14 cubic inches per minute, how long will it take for all the liquid in the funnel to pass through the nozzle? (Use π = 3.14.) (1 point) 4.71 minutes 3.14 minutes 14.13 minutes 9.42 minutes

OpenStudy (anonymous):

How do you upload an image?

OpenStudy (mertsj):

We're talking about a cone here so just calculate the volume of the cone in cubic inches.

OpenStudy (mertsj):

Divide that answer by 14 and that will be the number of minutes.

OpenStudy (anonymous):

The image is a conical funnel

OpenStudy (mertsj):

Find the volume of the cone.

OpenStudy (anonymous):

the volume of the cone is 9.42, right?

OpenStudy (mertsj):

What is the radius of the cone?

OpenStudy (anonymous):

3 in

OpenStudy (mertsj):

What is the height of the cone?

OpenStudy (anonymous):

oh wait I forgot to multiply the height!

OpenStudy (anonymous):

the height is 7 in

OpenStudy (anonymous):

65.94 is the volume

OpenStudy (anonymous):

I think the answer is 4.71 correct?

OpenStudy (mertsj):

\[V=\frac{1}{3}(3.14)(3^2)(7)=65.94\in^3\]

OpenStudy (anonymous):

ok that's what I got

OpenStudy (mertsj):

\[\frac{65.94}{14}=4.71minutes\]

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

Do you mind helping me with some more? I don't what to keep you though.